express y in terms of x:
a)log y= log x+2
b)ln y= ln 2-3 ln x
c)log y= 2.5x +1

Answer by Nate(3500)

a)log y = log x+2
y = x + 2
b)ln y = ln 2 - 3ln x
ln y = ln 2 - ln x^3
ln y = ln 2/x^3
y = 2/(x^3)
c)log y = 2.5x +1
y = 10^(2.5x + 1)